# Energy-Efficient Approximate Full Adders Applying Memristive Serial IMPLY Logic For Image Processingthanks: Preprint Submitted to arXiv

### v1

@@ Abstract
Researchers and designers are facing problems with memory and power walls, considering the pervasiveness of Von-Neumann architecture in the design of processors and the problems caused by reducing the dimensions of deep sub-micron transistors. Memristive Approximate Computing (AC) and In-Memory Processing (IMP) can be promising solutions to these problems. We have tried to solve power and memory wall problems by presenting the implementation algorithm of four memristive approximate full adders applying the Material Implication (IMPLY) method. The proposed circuits reduce the number of computational steps by up to 40% compared to State-of-the-art (SOA). The energy consumption of the proposed circuits improves over the previous exact ones by 49%-75% and over the approximate full adders by up to 41%. Multiple error evaluation criteria evaluate the computational accuracy of the proposed approximate full adders in three scenarios in the 8-bit approximate adder structure. The proposed approximate full adders are evaluated in three image processing applications in three scenarios. The results of application-level simulation indicate that the four proposed circuits can be applied in all three scenarios, considering the acceptable image quality metrics of the output images (the Peak Signal to Noise Ratio (PSNR) of the output images is greater than 30 dB).
@@ 1 Introduction
The power wall is a considerable adversity in the computer architects’ path [1]. Energy consumption in portable electronic devices becomes a significant challenge with the invalidation of Dennard’s scaling, deviation from Moore’s law, and the occurrence of problems with the reduction of transistor dimensions, such as increased leakage currents, short channel effects, reduced reliability, and increased manufacturing costs [2, 3, 4, 5, 6, 7, 8]. The difference in the speed of memories and processors and Von-Neumann’s memory wall bottleneck is another challenge in the design of digital systems [1, 9, 10, 11].
@@
Several solutions to overcome the power wall problem have been proposed to date. One of the solutions welcomed by researchers and technology companies like IBM in recent years is AC [2, 3, 4, 12, 13, 14]. In AC, the accuracy of computations is reduced to an acceptable level. In exchange for this reduction in accuracy, designers can reduce circuit complexity (energy consumption, delay, and area). AC cannot be applied in all aspects of computations because the accuracy of computations must be 100% in some processing applications. Error-resistant applications such as image processing, pattern recognition, machine learning, digital signal processing, and data mining are among the applications in which AC can be applied [2, 4, 14, 15, 16, 17]. For example, in image processing, reasons such as a small error in the output pixels due to data noise or approximately performing computations affect the quality of the images [2, 15, 17]. However, humans’ limited visual perception accepts the results [2, 3, 4, 17, 18].
@@
The idea of applying processing units next to the main memory (due to the impossibility of data processing by the Dynamic Random Access Memory (DRAM) cell) has been applied by researchers for decades to overcome the memory wall. One of the proposed methods to overcome the memory wall is IMP. With the advancement of technology and the introduction of memristors, it is possible to apply IMP more than ever before. A memristor is an electrical element that maintains its previous resistance value when voltage and current are not applied to it and is contemplated as a memory cell [19, 20, 21]. As a memory element and in stateful logic, the High Resistance State (HRS) of a memristor is considered as logic ‘0’, and its Low Resistance State (LRS) is considered as logic ‘1’ [3, 10, 14, 22]. Different memory technologies have been contemplated for IMP. However, Resistive RAM (ReRAM) is a popular option among emerging memory technologies for IMP due to its small size, suitable writing speed, low energy consumption, and acceptable endurance [11, 19, 23, 24]. ReRAMs fit well in crossbar array structure and harmonize entirely with current manufacturing technology [3, 10, 19]. There are several methods for designing computing structures applying memristors. Among the most important of these methods, Memristor Aided loGIC (MAGIC), fast and energy-efficient logic in memory (FELIX), and IMPLY can be mentioned [11, 19, 21, 24, 25, 26].
@@
Arithmetic circuits are one of the most critical units of each processor. Adders are the central pillar of computing circuits. For example, many instructions in digital signal processing (DSP) processors require adder and multiplier units [3, 27]. So far, many approximate adders have been introduced in different technologies for error-resilient applications that fit today’s typical architecture [3, 4, 13, 14, 15, 16, 18, 28, 29, 30]. In [3, 14, 31, 32], approximate memristor-based full adders are proposed by applying the IMPLY logic for IMP. In this article, AC and IMP by memristors and IMPLY method are applied side by side, and four approximate full adders are proposed for error-resilient applications. Contributions made in this article include:
@@
    Proposing a compact approximate full adder with Exact $C_{out}$ and Inexact $Sum$ (ECIS) to prevent the propagation of the wrong carry digit from the Least Significant Bits (LSBs) to the Most Significant Bits (MSBs) and improve energy consumption and the number of computational steps compared to the exact full adders,
@@
    Presenting a Figure of Merit ($FOM$) to evaluate the proposed circuits by circuit evaluation criteria and accuracy metrics simultaneously.
The rest of the article is arranged in 4 sections. In the second section, the backgrounds of this research are explained, and SOA is briefly introduced and reviewed. In the third section, the implementation algorithms of ECIS and ICIS1-3 full adders are introduced, and their design details are explained. The fourth section evaluates and compares the proposed circuits and SOA by circuit evaluation criteria, error analysis metrics, and image quality metrics. In subsection 4.4, the proposed approximate full adders are also evaluated by a $FOM$ to have a comprehensive view of the reduction of circuit complexity and the reduction of accuracy in the computations of the proposed circuits. Section 5 contains the conclusion of the article.
2 Backgrounds
@@ 2.1 Approximate computing
Reducing the accuracy of computations by applying AC in data-intensive error-resilient applications, such as image processing, can unravel the hardware complexity of these applications. AC has been applied to design circuits from memory to arithmetic circuits, such as adders and multipliers [3, 4, 5, 13, 14, 15, 16, 17, 18, 27, 28, 29, 30]. Designers should focus on the compromise between reducing the accuracy of computations and hardware complexity (reducing energy consumption and increasing efficiency) in applications where AC is applied. The reasons for applying AC can be limited visual perception of humans, lack of only one acceptable answer, resistance to input noise, and error coverage and attenuation [2, 17]. Among the essential error-resilient applications are image processing, deep learning, robotics, data mining, and pattern recognition [2, 15, 16, 17].
To design approximate arithmetic circuits, circuit evaluation criteria such as energy consumption, performance, and area dissipation should be evaluated to ensure the improvement of circuit evaluation criteria. The evaluation of the accuracy of computations is critical from two different aspects. The first aspect is the examination of standard error evaluation criteria such as Error Rate (ER), Error Distance (ED), Mean ED (MED), and Normalized MED [15]. By examining these criteria, one can get a general view of the accuracy of computations by applying the designed approximate circuit in different computing structures. The method of calculating these error evaluation criteria is described in [3, 4, 15, 16]. In general, the smaller the error evaluation criteria, the higher the accuracy of the computation. The second aspect is to pay attention to the application in which the proposed circuits are applied. The specific output evaluation criteria of the desired application should also evaluate the designed approximate arithmetic circuits. In the application of image processing, in which the proposed approximate circuits are applied in this paper, image quality evaluation criteria such as PSNR, Structural Similarity Index (SSIM), and Mean SSIM (MSSIM) should be assessed to ensure the proper performance of the approximate arithmetic circuits [3, 4, 12, 13, 14, 15, 16, 17, 18, 33]. The criteria for evaluating the quality of images and their importance were studied in [12, 34]. The greater the image quality evaluation criteria, the better the quality of the output images. The output image is acceptable and of good quality when the image created by the approximate circuit has a PSNR greater than 30 dB [33].
@@ 2.2 Memristors
Emerging non-volatile memory technologies have attracted the attention of researchers in recent years, considering the importance of memory cells in all processing structures and the limitations caused by reducing the dimensions of transistors, especially in flash memories [19]. A memristor is an electrical element introduced by Leon Chua in 1971 [19]. Memristors can keep their previous state, and this feature makes them a suitable alternative to today’s standard memory technologies. ReRAM has become a prevalent choice for replacing conventional memory technologies among researchers due to its unique features compared to other emerging memory technologies such as Ferroelectric RAM (FeRAM), Phase Change Memory (PCM), and Magnetic RAM (MRAM) [19, 35].
So far, several logical design methods have been presented for circuit design applying memristors [11, 19, 21, 24, 26]. Each of these methods has its characteristics. Statefulness is one of the essential features to consider for implementing memristor-based circuits [3, 19, 25, 36]. In the implementation of a circuit, if the input, intermediate nodes, and output states of a circuit are represented by resistance, the design method of that circuit is stateful [19]. In order to be able to perform IMP applying memristors, stateful design methods should be applied [19, 25]. One of the stateful and flexible memristor-based circuit design methods introduced by Hewlett-Packard (HP) is IMPLY [3, 14, 19]. The structure of the IMPLY function is shown in Figure 1, and its truth table is written in Table 1. In IMPLY logic, the inputs are set to HRS and LRS in memristors $p$ and $q$, and after performing the IMPLY operation, the result is written in memristor $q$ [21]. The following conditions must be met to perform IMPLY correctly [3, 10, 14, 21, 22]:
@@
    $V_{COND}<V_{c}<V_{SET}$, $V_{c}$ is the threshold voltage of a memristor.
@@
Figure 1: Circuit design of a memristive IMPLY logic gate [10].
@@ 2.3.1 IMPLY-based adders’ architectures
Four serial, parallel, semi-serial, and semi-parallel architectures were introduced to implement the algorithms of arithmetic structures (e.g., n-bit adders) applying IMPLY method [10, 11, 20, 21, 22].
In serial architecture, memristors are placed next to each other in a row/column and connected to the ground by a resistor. Only a FALSE or an IMPLY operation can be executed in each computational step. The serial architecture of an n-bit adder is shown in Figure 2. Serial architecture has the simplest structure compared to other architectures. It can be well implemented in the structure of crossbar arrays [20]. Due to the simplicity of the serial architecture and the impossibility of executing two instructions simultaneously, the number of computational steps in this method is more than in the other architectures.
In parallel architecture, memristors are placed in rows (or columns) parallel to each other. Each row/column is connected to a resistor of that row/column by a switch and to other rows/columns by another switch (see Figure 3). If there is no dependency between the memristors of consecutive steps in an implementation algorithm, the instructions applying this architecture can be calculated simultaneously [37]. However, suppose there is a dependency between the memristors of consecutive instructions. In that case, it is impossible to execute several instructions in parallel, e.g., to calculate the $C_{out}$ of a full adder proposed in [37]. Parallel architecture has the maximum hardware complexity compared to other architectures, but the number of computational steps of different algorithms in this architecture is less than in the other architectures [10, 22].
In [10, 22], semi-parallel and semi-serial architectures for implementing n-bit adders are introduced. In both architectures, the input memristors (input numbers $A$ and $B$) are placed in two separate rows. In the semi-parallel architecture (see Figure 4), there is a work memristor in each line ($w_{1}$ and $w_{2}$), and the $C_{in}$ memristor is placed in the second row [22]. Each row is connected to a resistor by a switch, and another switch connects two rows [22]. In the semi-serial method (see Figure 5), each row is directly connected to a resistor [10]. Each of the work memristors ($c$ and $w_{1}$-$w_{4}$), along with the $C_{in}$ memristor, can be connected to the input memristors of each of the upper or lower rows by two different switches and perform the desired operation in this architecture [10].
Figure 2: The serial architecture of an IMPLY-based n-bit adder [10].
Figure 3: The parallel architecture of an IMPLY-based n-bit adder [10].
Figure 4: The semi-parallel architecture of an IMPLY-based n-bit adder [10].
Figure 5: The semi-serial architecture of an IMPLY-based n-bit adder [10].
@@ 2.3.2 IMPLY-based exact full adders
The importance of the full adder cell in all kinds of computing structures suitable for Von-Neumann architecture and IMP is obvious. For this reason, researchers have introduced multiple IMPLY-based full adders in recent years [10, 20, 21, 22, 37]. These full adders are designed based on all four architectures introduced in section 2.3.1. The number of required memristors, computational steps, and CMOS switches are among the most critical criteria differentiating these circuits. In Table 2, some of the most prominent IMPLY-based exact full adders introduced are compared with each other along with their features in the architecture applied [10, 20, 21, 22, 37].
@@ 2.3.3 IMPLY-based approximate full adders
To the best of our knowledge, the first approximate full adder based on the IMPLY logic was introduced in [14]. This full adder computes $Sum$ and $C_{out}$ outputs in 8 computational steps with only five memristors [14]. The main problem of this approximate full adder is that the outputs are written in work memristors. As the number of approximate full adders increases in the n-bit approximate adder structure, the number of required memristors increases. The authors in [3] solved the problem of their previously proposed approximate full adder introduced in [14] by modifying its implementation algorithm. In [3], four Serial IMPLY-based approximate full adders (SIAFA1-4) are introduced. SIAFA1-4 require four/five memristors (three input memristors and one/two work memristors) to implement and compute the outputs in 8, 10, 8, and 8 computational steps, respectively [3]. Serial Approximate Full Adder using NAND gates (SAFAN) is another IMPLY-based approximate full adder that was proposed recently [31]. This approximate full adder is designed and implemented in serial architecture like the ones proposed in [3, 14]. The implementation algorithm of SAFAN computes the outputs in seven computational steps by applying four memristors (three input memristors and one work memristor) [31]. Seiler et al. proposed an approximate full adder cell (semi-serial AFA) compatible with the semi-serial architecture [32]. The $Sum$ and $C_{out}$ outputs of this full adder are computed in six computational steps with five memristors [32]. Error evaluation metrics of ED, MED, and NMED were calculated for the approximate full adders proposed in [3, 31, 32]. The circuit and error evaluation criteria of SIAFA1-4, SAFAN, and semi-serial AFA are summarized and written in Table 3. SAFAN, SIAFA1-4, and semi-serial AFA was evaluated in image processing applications, too [3, 31, 32].
@@
2.3.4 An overview of non-stateful memristor-based approximate full adders
One of the non-stateful memristor-based arithmetic circuit design methods is the Memristor Ratioed Logic (MRL) [11, 19]. This method creates AND and OR gates by connecting two memristors in series according to their polarities [11, 19]. The logical states of the inputs and outputs of the MRL-based circuits are determined by voltage levels, like Complementary Metal Oxide Semiconductor (CMOS) circuits [11, 19]. There is a voltage drop problem in the MRL method’s output signals, so some intermediate nodes need buffers to restore the signal level [19]. Using these buffers and signal transmission between memristors, vias, and transistors is associated with energy loss and affects the delay [3].
Different exact and approximate full adders were introduced by applying this method [11, 19, 38, 39, 40]. An MRL-based exact full adder was introduced in [38, 39]. This full adder consists of 33 memristors and some inverters [38, 39]. Two MRL-based approximate full adders were proposed based on redefining approximate logic from the exact logic method in [38, 39]. These approximate full adders are implemented with 10 and 14 memristors and by applying some inverters to restore signal levels [38, 39]. These approximate full adders were evaluated in the application of image addition, and the output results were assessed with PSNR [3, 38, 39].
2.4 Approximate full adders in CMOS and emerging technologies
In more than a decade and due to the importance of the full adder cell, several approximate full adders applying MOSFET transistors were introduced as today’s standard manufacturing technology and evaluated in various error-resilient applications [4, 15, 16, 18]. Several approximate full adders were introduced in emerging technologies such as Carbon Nano Tube (CNT) FET and Quantum Cellular Automata (QCA), and their efficiency was evaluated in error-tolerant applications [28, 29, 30]. In most cases, these approximate circuits are designed based on redesigning the approximate logic from the exact logic method by rearranging the circuits and their input and altering the exact full adder’s truth table to the approximate one [4, 18, 28, 29, 30].

### v2

@@ Abstract
Researchers and designers are facing problems with memory and power walls, considering the pervasiveness of Von-Neumann architecture in the design of processors and the problems caused by reducing the dimensions of deep sub-micron transistors. Memristive Approximate Computing (AC) and In-Memory Processing (IMP) can be promising solutions to these problems. We have tried to solve the power and memory wall problems by presenting the implementation algorithm of four memristive approximate full adders applying the Material Implication (IMPLY) method. The proposed circuits reduce the number of computational steps by up to 40% compared to the state-of-the-art. The energy consumption of the proposed circuits improves over the previous exact ones by 49%-75% and over the approximate full adders by up to 41%. Multiple error evaluation criteria evaluate the computational accuracy of the proposed approximate full adders in three scenarios in the 8-bit approximate adder structure. The proposed approximate full adders are evaluated in three image processing applications in three scenarios. The results of application-level simulation indicate that the four proposed circuits can be applied in all three scenarios, considering the acceptable image quality metrics of the output images.
@@ 1 Introduction
The power wall is a considerable adversity in the computer architects’ path [1]. Energy consumption in portable electronic devices becomes a significant challenge with the invalidation of Dennard’s scaling, deviation from Moore’s law, and the occurrence of problems with the reduction of transistor dimensions, such as increased leakage currents and short channel effects [2, 3, 4, 5, 6, 7, 8, 9]. The growing difference in the speed of memories (the rate of data that can be supplied to the processor from main memory) and processors known as Von-Neumann’s memory wall bottleneck is another challenge in the design of digital systems [1, 10, 11, 12].
@@
Several solutions to overcome the power wall problem have been proposed to date. One of the solutions welcomed by researchers and technology companies like IBM in recent years is Approximate Computing (AC) [2, 3, 4, 13, 14, 15]. In AC, the accuracy of computations is reduced to an acceptable level. Designers can reduce circuit complexity as a result of this accuracy reduction [16]. Error-resistant applications such as image processing and machine learning are among the applications in which AC can be applied [2, 4, 15, 17, 18, 19, 20]. For example, in image processing, reasons such as a small error in the output pixels due to data noise or approximately performing computations affect the quality of the images [2, 17, 19]. However, humans’ limited visual perception accepts the results [2, 3, 4, 19, 21].
@@
The idea of applying processing units next to the main memory has been applied by researchers for decades to overcome the memory wall. With the advancement of technology and the introduction of memristors, it is possible to apply In-Memory Processing (IMP) more than ever before. A memristor is an electrical element that maintains its previous resistance value when voltage and current are not applied to it and is contemplated as a memory cell [22, 23, 24]. As a memory element and in stateful logic, the High Resistance State (HRS) of a memristor is considered as logic ‘0’, and its Low Resistance State (LRS) is considered as logic ‘1’ [3, 11, 15, 25]. Different memory technologies have been contemplated for IMP. However, Resistive RAM (ReRAM) is a popular option among emerging memory technologies for IMP due to its small size, suitable writing speed, low energy consumption, and acceptable endurance [12, 22, 26, 27]. ReRAMs fit well in crossbar array structure and harmonize entirely with current manufacturing technology [3, 11, 22]. There are several methods for designing computing structures that apply memristors. Among the most important of these methods, Memristor Aided loGIC (MAGIC), fast and energy-efficient logic in memory (FELIX), and Material Implication (IMPLY) are attractive [12, 22, 24, 27, 28, 29].
@@
Arithmetic circuits are one of the most critical units of each processor. Adders are the central pillar of computing circuits. For example, many instructions in Digital Signal Processing (DSP) processors require adder and multiplier units [3, 30]. So far, many approximate adders have been introduced in different technologies for error-resilient applications that fit today’s typical architecture [3, 4, 14, 15, 17, 18, 21, 31, 32, 33]. In [3, 15, 34, 35], approximate memristor-based full adders are proposed by applying the IMPLY logic for IMP. In this article, AC and IMP are applied alongside by memristors and IMPLY method, and four approximate full adders are proposed for error-resilient applications. The contributions made in this article are as follows:
@@
    Proposing a compact approximate full adder with Exact CarryoutCarry~out ($C_{out}$) and Inexact $Sum$ (ECIS) to prevent the propagation of the wrong carry digit from the Least Significant Bits (LSBs) to the Most Significant Bits (MSBs) and improve energy consumption and the number of computational steps compared to the exact full adders,
@@
    Presenting two Figures of Merit ($FOMs$) to evaluate the proposed circuits by circuit evaluation criteria and accuracy metrics simultaneously.
The rest of the article is arranged in 4 sections. In the second section, the background of this research are explained, and state-of-the-art is briefly introduced and reviewed. In the third section, the implementation algorithms of ECIS and ICIS1-3 full adders are introduced, and their design details are explained. The fourth section evaluates and compares the proposed circuits and state-of-the-art by circuit evaluation criteria, error analysis metrics, and image quality metrics. In subsection 4.4, the proposed approximate full adders are also evaluated by two $FOMs$ to have a comprehensive view of the reduction of circuit complexity and the reduction of accuracy in the computations of the proposed circuits. Section 5 contains the conclusion of the article.
2 Background
@@ 2.1 Approximate computing
Reducing the accuracy of computations by applying AC in data-intensive error-resilient applications can diminish the hardware complexity of these applications. AC has been applied to design circuits from memory to arithmetic circuits, such as adders and multipliers [3, 4, 5, 14, 15, 17, 18, 19, 21, 30, 31, 32, 33]. Designers should focus on the compromise between reducing the accuracy of computations and reducing energy consumption and increasing efficiency in applications where AC is applied. The reasons for applying AC can be limited visual perception of humans, lack of only one acceptable answer, resistance to input noise, and error coverage and attenuation [2, 19].
To design approximate arithmetic circuits, circuit evaluation criteria such as energy consumption, performance, and area dissipation should be evaluated to ensure the improvement of circuit evaluation criteria. The evaluation of the accuracy of computations is critical from two different aspects. The first aspect is the examination of standard error evaluation criteria such as Error Rate (ER), Error Distance (ED), Mean ED (MED), and Normalized MED [17]. By examining these criteria (see (1)-(4)), one can get a general view of the accuracy of computations by applying the designed approximate circuit in different computing structures. In general, the smaller the values of the error evaluation criteria, the higher the accuracy of the computation. The second aspect is to pay attention to the application in which the proposed circuits are applied. The specific output evaluation criteria of the desired application should also evaluate the designed approximate arithmetic circuits. In the application of image processing, in which the proposed approximate circuits are applied in this paper, image quality evaluation criteria such as Peak Signal to Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Mean SSIM (MSSIM) should be assessed to ensure the proper performance of the approximate arithmetic circuits [3, 4, 13, 14, 15, 17, 18, 19, 21, 36]. The criteria for evaluating the quality of images and their importance were studied in [13, 37]. The greater the image quality evaluation criteria, the better the quality of the output images. The output image is acceptable and of good quality when the image created by the approximate circuit has a PSNR greater than 30 dB [36].
$ER=\frac{No.~of~approximate~states}{Total~no.~of~states}$ (1)
$ED=\mid A-E\mid,A:~approximate~value~\&~E:~exact~value$ (2)
$MED=\frac{\sum ED}{2^{n}},n:No.~of~bits$ (3)
$NMED=\frac{MED}{MAX},MAX:~maximum~output~value~of~an~exact~circuit$ (4)
@@ 2.2 Memristors
Emerging non-volatile memory technologies have attracted the attention of researchers in recent years, considering the importance of memory cells in all processing structures and the limitations caused by reducing the dimensions of transistors, especially in flash memories [22]. Memristors can keep their previous state, and this feature makes them a suitable alternative to today’s standard memory technologies. ReRAM has become a prevalent choice for replacing conventional memory technologies among researchers due to its unique features [22, 38].
So far, several logical design methods have been presented for circuit design applying memristors [12, 22, 24, 27, 29]. Each of these methods has its characteristics. Statefulness is one of the essential features to consider for implementing memristor-based circuits [3, 22, 28, 39]. In the implementation of a circuit, if the input, intermediate nodes, and output logic states of a circuit are represented by resistance values, the design method of that circuit is stateful [22]. In order to be able to perform IMP applying memristors, stateful design methods should be applied [22, 28]. One of the stateful and flexible memristor-based circuit design methods introduced by Hewlett-Packard (HP) is IMPLY [3, 15, 22]. The IMPLY function is the primary function in the memristive IMPLY design method, done between two memristors (e.g., memristors $p$ and $q$). Applying the IMPLY function between $p$ and $q$ is represented by $p\to q$, which equals $\overline{p}+q$ [23, 24]. The structure of the IMPLY logic gate is shown in Figure 1(a), where the memristors are connected to the ground via $R_{G}$, and its truth table is written in Table 1. In IMPLY logic, first, the inputs are set to HRS or LRS in memristors $p$ and $q$, and after performing the IMPLY operation, by applying voltage $V_{COND}$ to memristor $p$ and voltage $V_{SET}$ to $q$, simultaneously, the result is written in memristor $q$ [24]. The following conditions must be met to perform IMPLY correctly [3, 11, 15, 24, 25]:
@@
    $V_{COND}<V_{c}<V_{SET}$, $V_{c}$ is the threshold voltage of a memristor,
@@
Figure 1: (a) Circuit design of a memristive IMPLY logic gate [3], and (b) the serial architecture of an IMPLY-based n-bit adder [3].
@@ 2.3.1 IMPLY-based adders’ architectures
Four serial, parallel, semi-serial, and semi-parallel architectures were introduced to implement the algorithms of arithmetic structures (e.g., n-bit adders) applying the IMPLY method [11, 12, 23, 24, 25].
In serial architecture, memristors are placed next to each other in a row/column and connected to the ground by a resistor. Only a FALSE or an IMPLY operation can be executed in each computational step. The serial architecture of an n-bit adder is shown in Figure 1(b). Serial architecture has the simplest structure compared to other architectures. It can be well implemented in the structure of crossbar arrays [23]. Due to the simplicity of the serial architecture and the impossibility of executing two computational steps simultaneously, the number of computational steps in this method is more than in the other architectures.
@@ 2.3.2 IMPLY-based exact full adders
The importance of the full adder cell in all kinds of computing structures suitable for the Von-Neumann architecture and IMP is obvious. For this reason, researchers have introduced multiple IMPLY-based full adders in recent years [11, 23, 24, 25, 40]. These full adders are designed based on all four architectures introduced in subsection 2.3.1. The number of required memristors, computational steps, and CMOS switches are among the most critical criteria differentiating these circuits. In Table 2, some of the most prominent IMPLY-based exact full adders introduced are compared with each other along with their features in the architecture applied [11, 23, 24, 25, 40].
@@ 2.3.3 IMPLY-based approximate full adders
To the best of our knowledge, the first approximate full adder based on the IMPLY logic was introduced in [15]. This full adder computes $Sum$ and $C_{out}$ outputs in 8 computational steps with only five memristors [15]. The main problem of this approximate full adder is that the outputs are written in work memristors. As the number of approximate full adders increases in the n-bit approximate adder structure, the number of required memristors increases. The authors in [3] solved the problem of their previously proposed approximate full adder introduced in [15] by modifying its implementation algorithm. In [3], four Serial IMPLY-based Approximate Full Adders (SIAFA1-4) are introduced. SIAFA1-4 require four/five memristors (three input memristors and one/two work memristors) to implement and compute the outputs in 8, 10, 8, and 8 computational steps, respectively [3]. Serial Approximate Full Adder using NAND gates (SAFAN) is another IMPLY-based approximate full adder that was proposed recently [34]. This approximate full adder is designed and implemented in serial architecture like the ones proposed in [3, 15]. The implementation algorithm of SAFAN computes the outputs in seven computational steps by applying four memristors (three input memristors and one work memristor) [34]. Seiler et al. proposed an approximate full adder cell (semi-serial AFA) compatible with the semi-serial architecture [35]. The $Sum$ and $C_{out}$ outputs of this full adder are computed in six computational steps with five memristors [35]. Error evaluation metrics of ED, MED, and NMED were calculated for the approximate full adders proposed in [3, 34, 35]. The circuit and error evaluation criteria of SIAFA1-4, SAFAN, and semi-serial AFA are summarized and written in Table 3. SAFAN, SIAFA1-4, and semi-serial AFA were evaluated in image processing applications, too [3, 34, 35].